OpenStudy (anonymous):

2. Point B is between A and C. If AB = 6 cm and AC = 17 cm, what is BC? (Points : 3) 6 cm 11 cm 23 cm 46 cm

OpenStudy (anonymous):

BC=AC-AB or BC=17-6 or BC=11 So, your answers: 11cm

OpenStudy (anonymous):

11cm :) Subtract 17cm to 6cm

OpenStudy (mertsj):

Did you draw a picture?

OpenStudy (mertsj):